# New possibilities in identification of binary choice models with fixed effects

### v5

@@ Abstract
We study the identification of binary choice models with fixed effects. We provide a condition called sign saturation and show that this condition is sufficient for the identification of the model. In particular, we can guarantee identification even with bounded regressors. We also show that without this condition, the model is not identified unless the error distribution belongs to a small class. The same sign saturation condition is also essential for identifying the sign of treatment effects. A test is provided to check the sign saturation condition and can be implemented using existing algorithms for the maximum score estimator.
@@ 1 Introduction
This paper considers panel models with binary outcomes in the presence of fixed effects. These models are convenient in economic analysis as they allow for fairly general unobserved individual heterogeneity. The nonlinear nature of the binary choice models makes it difficult to eliminate the fixed effects by differencing. Since we view the fixed effects or their conditional distribution as a nuisance parameter, the identification of the parameter of interest becomes tricky. In this paper, we provide new results and insights on what drives the identification of the coefficients on the covariates and what this means for applied work.
We observe independent and identically distributed (i.i.d) copies of $(Y,X)$, where $Y=(Y_{0},Y_{1})\in\{0,1\}\times\{0,1\}$ and $X=(X_{0},X_{1})\in\mathcal{X}_{0}\times\mathcal{X}_{1}$. The data generating process is
$Y_{t}=\mathbf{1}\{X_{t}^{\prime}\beta+\alpha\geq u_{t}\}\qquad t\in\{0,1\},$ (1)
where the fixed effects are represented by the scalar variable $\alpha$. This is a semiparametric model as the distribution of $\alpha$ given $X$ is unrestricted.
@@
In this paper, we provide a more precise picture than Chamberlain, (2010) by showing that identification without parametric assumptions on the error distribution is quite possible even for bounded $X$. The main motivation of this paper is to explore new possibilities without functional-form assumptions. This seems an important issue on stability. If identification crucially hinges on the imposed functional form, the reliability of the analysis could be a concern. After all, if the model parameter is unidentified under every distribution function other than the logistic one, how much should we trust this model? Hence, it seems helpful to explore a more robust setting. In this direction, this paper considers the identification issue and leaves the estimation problem to future research. We make the following contributions.
First, we provide tight and simple identification conditions without parametric restrictions on the error distribution. We show that a sufficient condition for identification is what we refer to as sign saturation (Assumption 2), which states that $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)$ and $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)$ are both strictly positive. This condition can hold with bounded regressors $X$ and allow for multiple discrete regressors and interaction terms. Moreover, we show that this is also a necessary condition for identification unless the distribution of $u_{t}$ is in a special class. Therefore, although we know (from Chamberlain, (2010)) that for bounded regressors the identification fails in some regions of the parameter space, our results pinpoint these regions: the identification fails exactly in regions where the sign saturation fails. Therefore, whether the regressors are bounded is not what really drives the identification. The identifiability is more closely related to the sign saturation condition.
A key advantage of the proposed sign saturation condition is its simplicity in that it is stated in terms of observed variables only and does not involve the parameter that we are trying to identify. Hence, it can be used to answer the question of whether identification holds under the distribution of observed data, thereby making identification directly testable. This is different from the Manski, (1987)-type condition. For example, suppose that $K=2$, $W_{1}$ is bounded and the conditional distribution $W_{2}\mid W_{1}$ has support $\mathbb{R}$ almost surely. The condition in Manski, (1987) tells us that $\beta$ is identified if $\beta_{2}\neq 0$. However, it does not seem obvious how to check $\beta_{2}\neq 0$ from the data. In contrast, under the sign saturation condition, we simply check $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)$ and $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)$.
@@
Third, we show that this condition can be equivalently stated as moment conditions: $\sup_{q\in\mathbb{R}^{K}}\rho(q)>0>\inf_{q\in\mathbb{R}^{K}}\rho(q)$, where $\rho(q)=E[\mathbf{1}\{W^{\prime}q\geq 0\}(Y_{1}-Y_{0})]$. We propose a simple test for this that does not require any tuning parameters such as bandwidth or number of basis functions. For bounded regressors, the existing literature highlights the risk of identification failure: the identification possibly (not definitely) fails. This generic warning may have discouraged the application of fixed effect models in practice. The proposed test serves as a more accurate diagnosis. Using the proposed procedure, the applied researcher can simply check the identifiability for the specific application. It turns out that the computation is already familiar to applied researchers in this area as it only involves computing the usual maximum score estimator and its bootstrapping.
One important objective of our work is to provide robust tools for empirical research. Although conditional maximum likelihood or any other approach based on logistic errors is popular in applied work, we argue that this assumption can be problematic in applications. If so, it might be prudent to leave the error distribution nonparametric. Since estimation and inference based on maximum score estimates are already available (e.g., Manski, (1987) and Cattaneo et al., (2020)), we only need to make sure that we can reasonably assume the identification. This paper fills this very need: tight and testable conditions for identification. As an empirical illustration, we provide a test for the logistic assumption and apply it to the well-studied example of female labor participation. We find evidence against the logistic assumption. Using the proposed test for the sign saturation condition, we find evidence that the model parameter can be identified up to scale without any parametric assumption on the error distribution.
@@
Our work is closely related to the literature of binary response models with logistic errors. Rasch, (1960), Andersen, (1970) and Chamberlain, (1980) consider the estimation and inference in this case. The logistic link function is also singled out in Chamberlain, (2010) as the “nice” link function: the identification does not require unbounded regressors and the Fisher information bound is strictly positive. Recently, the work of Davezies et al., (2023) points out that if there are more than 2 time periods, then the nice link function can be extended to a generalized version of the logistic distribution. Logistic errors are also assumed in the study of identification in dynamic models, including recent works of Honoré and Weidner, (2020), Khan et al., (2020) among many others. This suggests that imposing a special class of parametric structures is a useful strategy of achieving identification and this special class seems to be the logistic functions. We provide new insight on this. We show that this special class also includes those such that $\dot{G}(\cdot)$ is periodic, where $\dot{G}(\cdot)$ is the derivative of $\ln\frac{F(\cdot)}{1-F(\cdot)}$ and $F(\cdot)$ is the distribution function of $u_{t}$. If $F(\cdot)$ is logistic, then $\dot{G}(\cdot)$ is a constant function, which is clearly periodic. However, we find that identification is possible for any periodic $\dot{G}(\cdot)$. Although Chamberlain, (2010) tells us that for non-logistic distributions, identification fails in an open neighborhood, we show that for any periodic $\dot{G}(\cdot)$, identification must hold in another open neighborhood. Therefore, it seems to us that the truly problematic distributions are those with non-periodic $\dot{G}(\cdot)$. Indeed, outside this special class of periodic $\dot{G}(\cdot)$, we show that identification is impossible at every point if the sign saturation condition is not satisfied. Thus, for generic distribution functions, sign saturation is a necessary condition for identification. We summarize the relation between identifiability and error distribution in Table 2.
@@
Our work is also closely related to works that do not assume logistic errors. First, some results in the literature allow for bounded covariates but completely rule out discrete variables. Assumption 3.3 of Shi et al., (2018) gives a sufficient condition for identification under bounded support. As commented in the paper, their assumption essentially requires all covariates to be continuous. Continuous distribution on all observed covariates are also required by Assumption 6 of Toth, (2017) and Assumption 4’ of Gao et al., (2023). There are also results that handle nonseparable models, including Hoderlein and White, (2012), Chernozhukov et al., (2015) and Chernozhukov et al., (2019). They rely on the derivative of $E(Y_{1}-Y_{0}\mid X)$ with respect to certain components of $X$ to identify structural parameters, but the nature of derivatives also means that this approach does not give identification results of the coefficient of a discrete variable. Here, we allow one binary regressor, which is important in empirical studies as many treatment variables are binary. We also extend it to the case of multiple discrete regressors in Section 3.2. Second, there is also a related model in the cross-sectional setting, where we observe $(Y,X)$ from $Y=\mathbf{1}\{X^{\prime}\beta+u\geq 0\}$ with $P(u\leq 0\mid X)=1/2$. For this model, Corollary 4.1 of Horowitz, (2009) proposes a condition for the identification of $\beta$. Under the normalization of $\beta=(1,\beta_{-1}^{\prime})^{\prime}$ and the corresponding partition $X=(X_{1},X_{-1}^{\prime})^{\prime}$, the key requirement is that there exist some small $\delta>0$ and some set $N_{\delta}$ with $P(X_{-1}\in N_{\delta})>0$ such that for almost every $x_{-1}\in N_{\delta}$, the density of $X_{1}+X_{-1}^{\prime}\beta_{-1}$ conditional on $X_{-1}=x_{-1}$ is everywhere positive on $[-\delta,\delta]$. One might wonder if a similar condition can be stated for the panel setting. Like Manski, (1987), this condition involves the unknown parameter $\beta$: to determine the identifiability of $\beta$ from the observed data, we need to check the density of $X^{\prime}\beta$ condition on $X_{k}$ for some $k$. As discussed above, the sign saturation condition is stated only in terms of observed data and is thus directly testable. The proposed test also has no tricky tuning parameters or curse of dimensionality that plague many nonparametric estimators, especially in multiple dimensions. Another technical issue is that statements of densities might not give the most general conditions; for example, some strictly increasing distribution functions have zero density almost everywhere, e.g., Takács, (1978). Perhaps most importantly, none of the aforementioned works show whether their conditions are necessary.
@@
Our necessity results complement those in Pakes and Porter, (2024). By Proposition 3 therein, $\arg\max_{\beta}E(Y_{1}-Y_{0})\cdot\mathbf{1}\{W^{\prime}\beta\geq 0\}$, the set of maximizers of the maximum score criterion function of Manski, (1987), is the sharp identification set when we only assume the stationarity of the distribution $u_{t}\mid(X,\alpha)$. Our results show that this sharp identified reduces to a singleton up to scale under the sign saturation condition. A natural question is the necessity of this condition, especially if we are willing to impose additional assumptions, namely i.i.d errors that are independent of $(X,\alpha)$. By classical results, we know that under logistic errors, $\arg\max_{\beta}E(Y_{1}-Y_{0})\cdot\mathbf{1}\{W^{\prime}\beta\geq 0\}$ is not the sharp identified set. Our results in Section 3.3 further complete the puzzle by characterizing point identification as the periodicity of $\dot{G}(\cdot)$.
2 An empirical illustration
In this section, we provide an empirical example to illustrate the main motivation of the paper: the logistic assumption on the error can be misspecified and the model parameter can be identified without parametric assumptions on the error term.
Consider an empirical study on women’s labor force participation. This problem has been studied by previous works, see Chamberlain, (1984), Chay and Hyslop, (1998), Fernández-Val, (2009) and Khan et al., (2020) among many others. To fix ideas, let $X_{i,t}=(D_{i,t},R_{i,t})^{\prime}\in\mathbb{R}^{2}$, where $D_{i,t}$ denotes the number of children under the age of 6 in household $i$ in year $t$ and $R_{i,t}$ denotes the log income of the husband in household $i$ in year $t$. In the data set, we have 9 years data from 1980 to 1988 with 1461 households. [^1] This specification is used in Example 15.5 of Wooldridge, (2010). We do not claim that this is the most general specification as there are definitely others in the literature, especially those with a lagged dependent variable or time dummy. Here, the goal is to use one specification to illustrate how to implement the test in practice. The data set is used by Fernández-Val, (2009) and can be downloaded from Fernández-Val, (2022).
We start with how to test the logistic assumption. Consider the model in (1) under the assumption that $u_{t}$ is independent across $t$ and independent of $(X,\alpha)$ and has a standard logistic distribution, i.e., $P(u_{t}\leq a)=[\exp(-a)+1]^{-1}$ for any $a\in\mathbb{R}$. Under this assumption, the parameter $\beta$ can be estimated consistently at the $\sqrt{n}$-rate, see e.g., Chamberlain, (1984). One natural question is whether this assumption is supported by the data. By simple calculations, we have
$P\left(Y_{0}=1,Y_{1}=0\mid X,\alpha\right)=\frac{1}{\left[1+\exp(-X_{0}^{\prime}\beta-\alpha)\right]\cdot\left[1+\exp(X_{1}^{\prime}\beta+\alpha)\right]}$
and
$P\left(Y_{0}=0,Y_{1}=1\mid X,\alpha\right)=\frac{1}{\left[1+\exp(X_{0}^{\prime}\beta+\alpha)\right]\cdot\left[1+\exp(-X_{1}^{\prime}\beta-\alpha)\right]}.$
The conditional maximum likelihood arises from this observation:
$\frac{P\left(Y_{0}=1,Y_{1}=0\mid X,\alpha\right)}{P\left(Y_{0}=0,Y_{1}=1\mid X,\alpha\right)}=\exp(-W^{\prime}\beta),$ (2)
where $W=X_{1}-X_{0}$. Since this ratio does not depend on $\alpha$, we can write down a logistic model when we consider a sub-sample with $Y_{0}\neq Y_{1}$:
$\frac{P\left(Y_{0}<Y_{1}\mid X,\alpha\right)}{P\left(Y_{0}\neq Y_{1}\mid X,\alpha\right)}=[\exp(-W^{\prime}\beta)+1]^{-1}.$
Therefore, we can simply restrict the sample to observations satisfying $Y_{0}\neq Y_{1}$ and simply run a logistic regression of $\mathbf{1}\{Y_{0}<Y_{1}\}$ on $W$ to identify $\beta$.
We now exploit this idea by rewriting (2) as the following conditional moment condition:
$P\left(Y_{0}=0,Y_{1}=1\mid X,\alpha\right)-P\left(Y_{0}=1,Y_{1}=0\mid X,\alpha\right)\cdot\exp(W^{\prime}\beta)=0,$
which implies
$E[\phi(Y,X;\beta)\mid X]=0,$ (3)
where $\phi(Y,X;\beta)=\mathbf{1}\{Y_{0}=0,Y_{1}=1\}-\mathbf{1}\{Y_{0}=1,Y_{1}=0\}\cdot\exp(W^{\prime}\beta)$. A simple way of checking (3) is to form the following unconditional moment condition:
${\rm there\ exists}\ \beta\ {\rm such\ that\ }E\left[\phi(Y,X;\beta)\cdot\begin{pmatrix}1\\ W\end{pmatrix}\right]=0.$ (4)
Notice that in (4), there are $\dim(\beta)+1$ moments. Hence, we can simply test the over-identifying conditions by the standard Hansen, (1982)’s $J$-test. [^2] Under the correct specification of the logistic distribution, we can consistently estimate the parameter by conditional maximum likelihood, which yields a consistent estimator for the optimal weighting matrix for the moment condition in (4). The limiting distribution of the $J$-statistic is the chi-squared distribution with one degree of freedom. The above strategy applies to the two-period case. Although there are 9 periods in our empirical example, we apply the above testing idea to each pair of periods. Instead of simultaneously testing the specification in all the time periods, this pairwise analysis allows us to visualize how often the logistic assumption holds. If the logistic assumption largely holds except a few specifications, then we can consider removing these problematic time periods; if we see widespread failure of the logistic assumption, then perhaps it is wise to abandon this assumption all together and seek alternative ways.
Table 1: P-value for $J$-test
$t=0\backslash t=1$ 1981 1982 1983 1984 1985 1986 1987 1988
1980 0. 3942 0. 2276 0. 9706 0. 0023 0. 0183 0. 0724 0. 4900 0. 5094
1981 0. 5776 0. 2130 0. 0001 0. 0045 0. 0221 0. 1980 0. 2602
1982 0. 0694 0. $0$ 0. $0$ 0. 0010 0. 0471 0. 0196
1983 0. 0009 0. 0088 0. 0989 0. 9220 0. 9741
1984 0. 4477 0. 4029 0. 0809 0. 0374
1985 0. 9964 0. 5536 0. 7115
1986 0. 5048 0. 9815
1987 0. 2135
We compute the p-value in the $J$-test for any pair of two years. For example, if we use 1980 and 1983 as the two time periods, then the p-value of the $J$-test for (4) is 0.3942. There are 36 pairs in total.
We report the p-values of the $J$-tests in Table 1. In 13 out of the (${9\choose 2}=36$ tests, the p-values are below 5%. Hence, this casts doubts on the validity of the logistic assumption in $13/36=36\%$ of all the two-period models. The logistic assumption is also rejected at the 5% level in a joint hypothesis testing; we see that there are several p-values below 0.05/36=0.0014, which is the nominal level after a simple Bonferroni-adjusted 5% test.
After seeing evidence against the logistic distribution assumption, we check whether or not $\beta$ can be identified without any parametric assumptions on the error distribution. As we shall show in Section 3, a sufficient and necessary condition for identification is the sign saturation condition. We now test $H_{0}$: the sign saturation condition fails versus $H_{1}$: the sign saturation condition holds. Under the test proposed in Section 5, the p-value is 0.0215 with 2000 bootstrap samples. Therefore, we have empirical evidence that the sign saturation condition holds and thus $\beta$ can be identified up to scale without relying on the logistic assumption.
3 Main results

### v7

@@ Abstract
We study the identification of binary choice models with fixed effects. We propose a condition called sign saturation and show that this condition is sufficient for identifying the model. In particular, this condition can guarantee identification even when all the regressors are bounded, including multiple discrete regressors. We also establish that without this condition, the model is not identified unless the error distribution belongs to a special class. Moreover, we show that sign saturation is also essential for identifying the sign of treatment effects. Finally, we introduce a measure for sign saturation and develop tools for its estimation and inference.
@@ 1 Introduction
This paper considers panel models with binary outcomes in the presence of fixed effects. These models are convenient in economic analysis as they allow for fairly general unobserved individual heterogeneity. The nonlinear nature of the binary choice models makes it difficult to eliminate the fixed effects by differencing. Since we view the fixed effects or their conditional distribution as a nuisance parameter, the identification of the parameter of interest becomes tricky. In this paper, we provide new results and insights on what drives the identification of the coefficients on the regressors and what this means for applied work.
Consider independent and identically distributed (i.i.d) observations $\{(Y_{i},X_{i})\}_{i=1}^{n}$ with $Y_{i}=(Y_{i,0},Y_{i,1})$ and $X_{i}=(X_{i,0},X_{i,1})$ from the following model
$Y_{i,t}=\mathbf{1}\{X_{i,t}^{\prime}\beta+\alpha_{i}\geq u_{i,t}\}\qquad t\in\{0,1\},$
where the fixed effects are represented by the scalar variable $\alpha_{i}$ and $\beta$ is a non-random vector of coefficients. This is a semiparametric model as the distribution of $\alpha_{i}$ given $X_{i}$ is unrestricted. For notational simplicity, we drop the $i$ subscript in the rest of the paper. Therefore, we write $Y=(Y_{0},Y_{1})$ and $X=(X_{0},X_{1})\in\mathcal{X}_{0}\times\mathcal{X}_{1}$ with
$Y_{t}=\mathbf{1}\{X_{t}^{\prime}\beta+\alpha\geq u_{t}\}\qquad t\in\{0,1\}.$ (1)
@@
In this paper, we provide a more precise picture than Chamberlain, (2010) by showing that identification without parametric assumptions on the error distribution is quite possible even for bounded $X$. The main motivation of this paper is to explore new possibilities without functional-form assumptions. This raises concerns about stability. If identification crucially hinges on the imposed functional form, the reliability of the analysis could be a concern. After all, if the model parameter is unidentified under every distribution function other than the logistic one, how much should we trust this model? Hence, it is helpful to explore a more robust setting. In this direction, this paper considers the identification issue and leaves the estimation problem to future research. We make the following contributions.
First, we provide tight and simple identification conditions without parametric restrictions on the error distribution. We show that a sufficient condition for identification is what we refer to as sign saturation (Assumption 2), which states that $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)$ and $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)$ are both strictly positive. This condition can hold with bounded regressors $X$ and allow for multiple discrete regressors and interaction terms. Moreover, we show that this is also a necessary condition for identification unless the distribution of $u_{t}$ is in a special class. Therefore, although we know (from Chamberlain, (2010)) that for bounded regressors the identification fails at some values of $\beta$, our results pinpoint these values: the identification fails exactly at points where the sign saturation fails. Therefore, whether the regressors are bounded is not what really drives the identification. The identifiability is more closely related to the sign saturation condition, which is simple and intuitive.
A key advantage of the proposed sign saturation condition is that it is stated in terms of observed variables only and does not involve the parameter that we are trying to identify. Hence, it can be used to answer the question of whether identification holds under the distribution of observed data, thereby making identification directly testable. This is different from the Manski, (1987)-type condition. For example, suppose that $K=2$, $W_{1}$ is bounded and the conditional distribution $W_{2}\mid W_{1}$ has support $\mathbb{R}$ almost surely. The identification condition in Manski, (1987) becomes $\beta_{2}\neq 0$. However, how can we check $\beta_{2}\neq 0$ when the identifiability of $\beta$ is not yet established? It does not seem obvious how to do so with the data. In contrast, under the sign saturation condition, we check statements only on observed data: $P\left(E(Y_{1}-Y_{0}\mid X)>0\right)$ and $P\left(E(Y_{1}-Y_{0}\mid X)<0\right)$.
@@
Third, we introduce a tool that can be used to assess the sign saturation condition empirically. We show that $E\max\{P(Y_{1}-Y_{0}\mid X),0\}$ and $E\min\{P(Y_{1}-Y_{0}\mid X),0\}$ can be directly estimated from the data and confidence intervals can be constructed by a simple bootstrap without any tuning parameters such as bandwidth or number of basis functions. For bounded regressors, the existing literature highlights the risk of identification failure: the identification possibly (not definitely) fails. This generic warning may have discouraged the application of fixed effect models in practice. The proposed tool serves as a more accurate diagnosis so that the identifiability can be assessed from the data.
@@
Our work is closely related to the literature of binary response models with logistic errors. Rasch, (1960), Andersen, (1970) and Chamberlain, (1980) consider the estimation and inference in this case. The logistic link function is also singled out in Chamberlain, (2010) as the “nice” link function: the identification does not require unbounded regressors. Recently, the work of Davezies et al., (2023) points out that if there are more than 2 time periods, then the nice link function can be extended to a generalized version of the logistic distribution. This suggests that imposing a special class of parametric structures is a useful strategy of achieving identification and this special class seems to be the logistic functions. We provide new insight on this. We show that this special class also includes those such that $\dot{G}(\cdot)$ is periodic, where $\dot{G}(\cdot)$ is the derivative of $\ln\frac{F(\cdot)}{1-F(\cdot)}$ and $F(\cdot)$ is the distribution function of $u_{t}$. If $F(\cdot)$ is logistic, then $\dot{G}(\cdot)$ is a constant function, which is clearly periodic. However, we find that identification is possible for any periodic $\dot{G}(\cdot)$. Although Chamberlain, (2010) tells us that for non-logistic distributions, identification fails in an open neighborhood, we show that for any periodic $\dot{G}(\cdot)$, identification must hold in another open neighborhood. Therefore, it seems to us that the truly problematic distributions are those with non-periodic $\dot{G}(\cdot)$. Indeed, outside this special class of periodic $\dot{G}(\cdot)$, we show that identification is impossible at every point if the sign saturation condition is not satisfied. Thus, for generic distribution functions, sign saturation is a necessary condition for identification. We summarize the relation between identifiability and error distribution in Table 1.
@@
Our work is also closely related to works that do not assume logistic errors. First, some results in the literature allow for bounded regressors but require all regressors to be continuous. For example, Assumption 3.3 of Shi et al., (2018) gives a sufficient condition for identification under bounded support. As commented in the paper, their assumption essentially requires all regressors to be continuous. Continuous distribution on all the regressors are also required by Assumption 6 of Toth, (2017) and Assumption 4’ of Gao et al., (2023). Examples on nonseparable models include Hoderlein and White, (2012), Chernozhukov et al., (2015) and Chernozhukov et al., (2019); their arguments rely on the derivatives with respect to $X$ and thus $X$ needs to be continuous. Here, we allow for one or multiple discrete regressors, which are important in empirical studies as many treatment variables are binary. Second, Corollary 4.1 of Horowitz, (2009) considers a related model in the cross-sectional setting. It is possible to translate this to the panel data in terms of conditional densities, but these are still not the most general conditions; for example, some strictly increasing continuous distribution functions have zero density almost everywhere, e.g., Salem, (1943) and Takács, (1978). Most importantly, none of the aforementioned works show whether their conditions are necessary. We contribute to the literature by finding what must be assumed and developing empirical tools to assess it.
@@
Our necessity results complement those in Pakes and Porter, (2024). By Proposition 3 therein, $\arg\max_{\beta}E(Y_{1}-Y_{0})\cdot\mathbf{1}\{W^{\prime}\beta\geq 0\}$, the set of maximizers of the maximum score criterion function, is the sharp identification set when we only assume the stationarity of the distribution $u_{t}\mid(X,\alpha)$. Our results show that this sharp identified reduces to a singleton up to scale under the sign saturation condition. A natural question is the necessity of this condition, especially if we are willing to impose additional assumptions, namely i.i.d errors that are independent of $(X,\alpha)$. By classical results, we know that under logistic errors, $\arg\max_{\beta}E(Y_{1}-Y_{0})\cdot\mathbf{1}\{W^{\prime}\beta\geq 0\}$ is not the sharp identified set. Our results in Section 2.3 further complete the puzzle by characterizing point identification as the periodicity of $\dot{G}(\cdot)$.
2 Identification of coefficients